OpenStudy (anonymous):

find the real square root . ( i just need an explanation on how to find it) 1.400

OpenStudy (loser66):

I give you another example to easily follow, since 400 has 2 zeros, it's not clear now find \(\sqrt{625}\) 

OpenStudy (loser66):

for first 6, like a division, which number whose square is lesser than 6? it s 2, because if it is 3, then 3^2 =9 >6, so, we pick 2, and 2^2 =4